OpenStudy (anonymous):

Four cards are drawn from a standard deck. No cards are returned. a) Which is more likely: • Event A: drawing a straight (four consecutive cards e.g. 5, 6, 7, 8)? • Event B: drawing four consecutive Queens? b) Explain how does your answer change if cards are returned after each draw?

OpenStudy (anonymous):

Event A is more likely...Pr(A)=(16/52)(15/51)(14/50)(13/49) .... and Pr(B)= (4/52)(3/51)(2/50)(1/49)

OpenStudy (anonymous):

if cards are returned then Pr(A)= (16/52)^4 ... and Pr(B)=(4/52)^4